3/5m+a=b, for m
please help, i have no i dea how 2 do these with fractions

Answer by tutorcecilia(2152)

3/5m+a=b
3/5m+a-a=b-a
3/5m=b-a
(5/3)3/5m=5/3(b-a)
m=5/3(b-a)
.
Check by plugging [m= 5/3(b-a) ] back into the original equation.
3/5m+a=b
3/5[5/3(b-a)]+a=b
b-a+a=b
b=b [Checks out]